OpenStudy (anonymous):

Which of the following is equivalent to this expression? 21a + 7 a (21 + 1) 7 (3a + 1) 7 (3a + 7) 3 (7a + 4)

OpenStudy (anonymous):

Medal given

OpenStudy (anonymous):

PLZ HELP

OpenStudy (tukitw):

You know how to factorize?

OpenStudy (anonymous):

Whats that?

OpenStudy (anonymous):

And no i dont

OpenStudy (anonymous):

@tukitw

OpenStudy (anonymous):

Will fan

OpenStudy (tukitw):

21a + 7 Notice that both 21 and 7 are divisible by 7, that is, 21 = 7 * 3 and 7 = 7 * 1. Since both numbers can be divided by 7, the common factor of the two terms would be 7. To factorise, we will take out the common factor. 21a + 7 = 7(3a + 1) The reason for the numbers inside the brackets is that 3 = 21 / 7 and 1 = 7 / 7.

OpenStudy (anonymous):

oh interesting

OpenStudy (anonymous):

Can u help me more